# UAV-borne Mapping Algorithms for Canopy-Level and High-Speed Drone Applications

### v1

@@ 2.2.3 DSOL: Direct Sparse Odometry Lite
DSOL presents an enhanced version of DSO and SDSO, proposing several algorithmic and implementation improvements to significantly speed up computation. Key aspects of the methodology include: (1) utilizing an inverse compositional alignment method for frame tracking, improving accuracy and speed. (2) adapting a better stereo photometric bundle adjustment formulation compared to SDSO. (3) Simplifying keyframe creation and removal criteria from DSO, allowing for better utilization of computational resources and parallel processing. (4) Implementing algorithmic enhancements to streamline the computation process, making it more suitable for real-time applications, especially in resource-constrained environments. Compared to DSO and SDSO, the focus of DSOL is on mapping speed and efficiency while maintaining accuracy.
@@ 3 Methodology
This section initiates by exploring optimal practices for UAV systems operating in both daytime and nighttime settings, offering a theoretical basis for the selection of sensors. Subsequently, the section delves into detailing the data and metrics employed to assess the efficacy of mapping algorithms in drone applications. Synthetic data was prioritized over real-world data due to its accessibility and the presence of ground truth 3D mapping information. The evaluation primarily centers on appraising the geometric accuracy of the reconstructed map.
@@ 3.1 Sensors for UAV Mapping
  • LiDAR (Light Distance and Ranging) Sensors
  • Event Cameras
  • Conventional EO and IR Cameras
Our assessment considered leading examples of each sensor that would be potentially appropriate for integration and commercially available. The specifications of the sensors were then reviewed in terms of their ability to provide measurements that meet the requirements of UAV mapping. Based on this analysis, a determination was reached regarding the suitability of each sensor.
@@ 3.1.1 LiDAR
Figure 3: Several LiDAR sensors were evaluated for inclusion on the platform. Left to right are shown the Ouster OS1, the HRL131, the RIEGL miniVUX-HA, and the L3 Harris Tactical Geiger-Mode LiDAR sensors.
Fig. 3 shows several LiDAR sensors evaluated for inclusion in the platform payload. However, it was quickly determined that these devices would not be appropriate for the UAV mapping application. The shortcomings of these sensors are described in the list below:
@@
  • Measurement speed: LiDAR sensors typically scan at low rates (10-20 Hz) which makes the capture of a complete 3D scene geometry impractical for the rates required by high-speed flight.
@@ 3.1.2 Event Cameras
Fig. 4 shows several event camera sensors evaluated for inclusion in the platform payload. The key attractive aspect of event cameras that has sparked considerable interest from researchers and industry alike is the extremely high temporal accuracy. Specifically, event cameras can resolve intensity changes in the perceptual field at a temporal resolution of approximately $1\mu s$. For this reason, event cameras have been used in high-speed contexts.
@@
  • Resolution: Resolution is a key parameter for depth accuracy as discussed in the stereo reconstruction Sec. 2.1.2. Accuracy strongly ties to both resolution and pixel size, $\delta$, as shown in Fig. 0(b) and Fig. 2. Event camera resolution, 0.3 pixels, is a factor of 5-10 times lower than conventional image sensors, and the pixel size of $\delta=18\mu m$ is a factor of 6-18 times larger than conventional image sensors, e.g., the Sony IMX472 sensor has a resolution of 21 Mpixels and a pixel size of $3.3\mu m$.
@@
  • Latency: While the temporal resolution of event cameras is an impressive $1\mu s$, the latency of the measurements is on the order of $<1ms$. This latency is similar to that of high frame rate conventional image sensors with frame rates of +100 fps and similar $<1ms$ latency.
@@
Figure 4: A collection of event cameras commercially available from iniVation Corp.
@@
3.1.3 Conventional EO and IR Cameras
Conventional image sensors have many beneficial attributes that often make them the sensor of choice for perception designs that must satisfy low Size, Weight, and Power (SWaP) requirements. They are well-suited for UAV mapping tasks for several reasons:
[…]
@@ 3.1.3 Conventional EO and IR Cameras
  • Integration with Other Technologies: Cameras can be integrated with other sensors and technologies, such as Inertial Measurement Units (IMUs) and Global Positioning Systems (GPSs), to enhance their capabilities and improve accuracy.
@@
  • Daytime and Nighttime Versatility: Camera sensors are capable of sensing in both daylight and nighttime conditions. If high sensitivity is needed in both scenarios, it is possible to replace daytime image sensors with infrared image sensors during nighttime conditions. This can be done with minor modifications to the underlying software and algorithms.
@@
Conventional image sensors have many beneficial attributes that make these sensors attractive for use as part of the Bumblebee perception payload. These sensors have low SWaP requirements and can record ¿16M measurements at a time from the environment. These sensors can be combined with lens components that provide both wide-angle viewpoints, e.g., a ${}^{\circ}$ FOV via the fisheye lens, for omnidirectional perception and confined viewpoints, e.g., ${}^{\circ}$ FOV “standard” lens, for high fidelity target tracking and mapping. The intensive work required to integrate these sensors and develop optimized algorithms to process their data to work using onboard computing resources can be reused between daytime (EO) and nighttime (IR) sensing contexts.
In the realm of image sensors, those designed for both infrared (IR) and visible light often share common image processing algorithms, including basic processes like filtering, noise reduction, contrast enhancement, and image registration. Additionally, object detection, recognition, feature extraction, and image fusion algorithms can typically be adapted for both IR and visible light images, leveraging shared features and patterns. However, notable differences emerge, primarily related to spectral characteristics, illumination, noise, calibration, temperature considerations, environmental conditions, and the unique sensitivities of IR images to object materials. These distinctions necessitate adjustments in algorithms to address variations in contrast, object recognition, and material discrimination, showcasing the need for specialized approaches in certain contexts.
@@ 3.2.1 Environment Simulation
AirSim [22] was used to simulate the dynamics of the drone. AirSim, developed by Microsoft, stands as a groundbreaking and influential simulator that has become a cornerstone in the development of autonomous drones and robotics. What sets AirSim apart is its capacity to simulate complex and dynamic environments with exceptional fidelity, replicating not only the physics of flight but also the intricacies of various sensors like cameras (RGB and depth), LiDAR, and GPS. Fig. 5 is an example of the AirSim simulated images captured by the cameras mounted on the multirotor.
Figure 5: An example of an AirSim environment showing the FPV view (main image) and simulated images at the bottom for cameras including stereo RGB (left and middle) and depth (right).
Although AirSim provides rich virtual environments for testing and fine-tuning a wide array of autonomous systems, these environments are often designed for games and lack realism. To simulate real-world flights and ultimately enhance the authenticity and effectiveness of simulations, this article used the Cesium plugin for the Unreal Engine, also known as “Unreal Cesium”. This plugin creates a powerful combination by integrating the Unreal Engine’s advanced rendering and simulation capabilities with Cesium’s geospatial visualization and data streaming features. By streaming high-resolution 3D models from Google Maps (latitude and longitude location required), overlying them onto real-world maps, and applying dynamic lighting and shadows, to provide precise representations of real-world locations, such as cities, terrains, and 3D models of buildings. This integration allows developers to create highly realistic and geospatially accurate virtual environments for various applications. Once the 3D map is generated, it behaves as a collision object in the Unreal Engine. Unmanned Aerial System (UAS) then interacts with this model using the geometry of the environment and a physics engine. Fig. 6 demonstrates the benefits of the Unreal Cesium plugin for simulation. It shows three real-world locations: (1) the UNC Charlotte campus, USA, (2) the Grand Canyon, USA, and (3) Paris, France.
(a) UNC Charlotte, NC, USA
(b) Grand Canyon, AZ, USA
(c) Eiffel Tower, Paris, France
Figure 6: Realistic environments created using Unreal Engine and Cesium Plug-in.
@@ 3.2.2 Flight Simulation
Fig. 7 depicts the proposed pipeline for flight simulation. QGroundControl and PX4-Autopilot are software components commonly used in the field of UAVs and drones. They work together to provide a comprehensive solution for controlling and managing drone flights. The missions are planned by defining waypoints, flight paths, and specific actions for the drone to perform and QGroundControl sends the mission plans to the autopilot system PX4-Autopilot. As an open-source flight control software for UAVs, PX4-Autopilot runs on the flight controller onboard the drone and is responsible for stabilizing the aircraft, executing flight plans, and interfacing with sensors and actuators. Controlled by PX4-Autopilot, the simulated UAV in AirSim follows the planned trajectory in a high-realism virtual environment created by Unreal Engine and Cesium plug-in. The cameras mounted on the UAV then capture the images (RGB, depth, etc) of the scene. These images along with the ground truth vehicle odometry can be obtained from AirSim, which can be used to reconstruct the ground truth 3D model of the world.
Figure 7: The flight simulation pipeline integrates 4 robot development technologies to facilitate development and testing: (1) Unreal Engine and Cesium Plug-in (high-realism image synthesis), (2) AirSim (vehicle dynamics), (2) QGroundControl (mission planning), and (4) PX4-Autopilot (vehicle control and Software-In-The-Loop).
@@ 3.3 Evaluation Methods
The evaluation of mapping performance encompasses two key dimensions: mapping accuracy and mapping speed. Mapping accuracy is meticulously assessed by comparing the geometry of the reconstructed point cloud with the ground truth point cloud. The ground truth point cloud is generated utilizing the accurate odometry and depth data provided in the simulated environment. Prior to the comparison process, the Iterative Closest Point (ICP) algorithm [1] is employed to establish point correspondences, associating points in the reconstructed point cloud with their counterparts in the ground truth cloud. Geometric accuracy serves as a pivotal metric, offering profound insights into the mapping algorithm’s ability to faithfully capture spatial relationships between reconstructed points and their true locations. Conversely, mapping speed quantifies the computational resources required for tracking frames and generating keyframes, which ultimately influences the efficiency of the reconstructed point cloud generation process. This comprehensive evaluation framework contributes to a nuanced understanding of the mapping algorithm’s overall performance, facilitating robust comparisons among different mapping methodologies.
3.3.1 Iterative Closest Point (ICP)
The Iterative Closest Point (ICP) algorithm [1] stands as a widely employed iterative optimization method for aligning two point clouds. Its primary objective is to determine the optimal transformation (comprising rotation, translation, and scale factor) that minimizes the distances between corresponding points in the two point clouds. Commencing with an initial estimate for the transformation, encompassing a scale factor $\mathbf{s}$, a rotation matrix $\mathbf{R}$, and a translation vector $\mathbf{t}$, the algorithm iterates through a sequence of steps. Initially, it establishes point correspondences between the source and target clouds by associating each point in the source with its nearest counterpart in the target. Subsequently, it calculates an error metric by evaluating the squared Euclidean distances among corresponding points. The algorithm then assesses convergence, and if the error falls below a predefined threshold, it terminates; otherwise, it proceeds to refine the transformation to further minimize the error. This iterative refinement process persists until either convergence is achieved or a predetermined maximum iteration count is reached.

### v2

@@ 2.2.3 DSOL: Direct Sparse Odometry Lite
DSOL presents an enhanced version of DSO and SDSO, proposing several algorithmic and implementation improvements to significantly speed up computation. Following the same practice as DSO of defining the photometric error in Equation (2), DSOL adopts the inverse compositional alignment method [baker2004lucas] to perform computationally expensive calculations, i.e., the Gauss-Newton approximation to the Hessian matrix, at the pre-computation phase, which largely improves the running speed of the algorithm. Compared to DSO and Stereo DSO, key aspects of optimization in DSOL include the following: (1) utilizing an inverse compositional alignment method for frame tracking, improving accuracy and speed; (2) adapting a better stereo photometric bundle adjustment formulation compared to SDSO; (3) simplifying keyframe creation and removal criteria from DSO, allowing for better utilization of computational resources and parallel processing; and (4) implementing algorithmic enhancements to streamline the computation process, making it more suitable for real-time applications, especially in resource-constrained environments. The focus of DSOL is on mapping speed and efficiency while maintaining accuracy.
2.3 Aerial Simulation Solutions
There are various simulation platforms for vehicles and environments catering to the diverse needs of researchers. Gazebo [gazebo_paper], with its open-source nature, stands as a versatile choice, emphasizing realism and adaptability. Agilicious [agilicious] specializes in agile quadrotor flight, providing unique applications such as drone racing. RotorS [rotors_sim], integrated with the Robot Operating System (ROS), offers high-fidelity UAV simulation. Flightmare [song2020flightmare], part of the AirSim project, excels in simulating multiple drones for swarm robotics research. Kumar Robotics Autonomous Flight [kumarRobotics1] addresses GPS-denied quadcopter autonomy. MIT’s FlightGoggles [flightgoggles] offers an immersive experience with photorealistic graphics. AirSim, developed by Microsoft, on top of the Unreal Engine, excels in generating highly realistic perceptual simulation data in complex and dynamic environments.
An approach is proposed in [beam2024cesium] to reproduce real-world experiments in simulation using the AirSim open-source simulator with the Cesium Tiles plugin, allowing for large-scale 3D geometry analysis. This paper adapts the methodology and extends it with other aerial vehicle control technologies, achieving precise vehicle control in high-realism virtual models that replicate real-world contexts world.
@@ 3 Methodology
The overall approach for the methods of this article consists of three steps:
  • Describe the benefits and shortcomings of various candidate sensing modalities for low-altitude high-speed mapping using Group 1 UAVs resulting in a recommendation for using one or more high-frame rate conventional camera sensors for this application (Section 3.1).
  • Describe the simulation methods used to collect data using a highly realistic 3D environment made possible by integrating the AirSim simulator with Google’s 3D map database using the Cesium Tiles plugin for the Unreal Engine (Section 3.2).
  • Describe the evaluation methods adopted to compare the mapping results generated from experimental flights within the simulated environment (Section 3.3).
@@ 3.1 Sensors for UAV Mapping
  • LiDAR (Light Distance and Ranging) Sensors;
  • Event Cameras;
  • Conventional EO and IR Cameras.
Our assessment considered leading examples of each sensor that would be potentially appropriate for the high-speed low-altitude context and commercially available. The specifications of the sensors were then reviewed in terms of their ability to provide measurements that meet the requirements of UAV mapping. Based on this analysis, a determination was reached regarding the suitability of each sensor.
@@ 3.1.1 LiDAR
Figure 3 shows several LiDAR sensors evaluated for inclusion in the platform payload. LiDAR sensors have emerged as a popular choice for UAV mapping applications with significant advancements in LiDAR-based techniques [li2019nrli, nguyen2021viral, lin2019evaluation, chiang2017development]. However, it was quickly determined that these devices would not be appropriate for the UAV mapping application. The shortcomings of these sensors are described in the list below:
@@
  • Measurement Speed: LiDAR sensors typically scan at low rates (10–20 Hz) which makes the capture of a complete 3D scene geometry impractical for the rates required by high-speed flight.
a
b
c
d
Figure 3: Several LiDAR sensors were evaluated for inclusion on the platform. Left to right are shown (a) the Ouster OS1, (b) the HRL131, (c) the RIEGL miniVUX-HA, and (d) the L3 Harris Tactical Geiger-Mode LiDAR sensors.
@@ 3.1.2 Event Cameras
Figure 4 shows several event camera sensors evaluated for inclusion in the platform payload. The key attractive aspect of event cameras that has sparked considerable interest from researchers and industry alike is the extremely high temporal accuracy. Specifically, event cameras can resolve intensity changes in the perceptual field at a temporal resolution of approximately 1 $\mu$s. For this reason, event cameras have been used in high-speed contexts.
Figure 4: A collection of event cameras commercially available from the iniVation Corp [eventcameras].
@@
  • Resolution: Resolution is a key parameter for depth accuracy as discussed in the stereo reconstruction Section 2.1.2. Accuracy strongly ties to both resolution and pixel size, $\delta$, as shown in Figures 1b and 2. Event camera resolution, 0.3 pixels, is a factor of 5–10 times lower than conventional image sensors, and the pixel size of $\delta$ = 18 $\mu$m is a factor of 6–18 times larger than conventional image sensors, e.g., the Sony IMX472 sensor has a resolution of 21 megapixels and a pixel size of 3.3 $\mu$m.
@@
  • Latency: While the temporal resolution of event cameras is an impressive 1 $\mu$s, the latency of the measurements is on the order of ¡1 ms. This latency is similar to that of high frame rate conventional image sensors with frame rates of +100 fps and similar ¡1 ms latency.
@@
3.1.3 Electro-Optical and Infrared Cameras
Conventional image sensors, including electro-optical (EO) and infrared (IR) sensors, have many beneficial attributes that often make them the sensor of choice for perception designs that must satisfy low Size, Weight, and Power (SWaP) requirements. They are well suited for UAV mapping tasks for several reasons:
  • SWaP: Both EO and IR camera modules are available commercially in a very large variety of form factors. This includes a compact 25 mm3 weighing 10–50 g requiring $\sim$1 W for power and providing temporally synchronized high framerate (60 fps) images.
@@ 3.1.3 Electro-Optical and Infrared Cameras
  • Integration with Other Technologies: Cameras can be integrated with other sensors and technologies, such as Inertial Measurement Units (IMUs) and Global Navigation Satellite System (GNSS), to enhance their capabilities and improve accuracy.
@@
  • Daytime and Nighttime Versatility: Camera sensors are capable of sensing in both daylight and nighttime conditions. If high sensitivity is needed in both scenarios, it is possible to replace daytime image sensors with infrared image sensors during nighttime conditions. This can be achieved with minor modifications to the underlying software and algorithms.
@@
Conventional image sensors have many beneficial attributes that make these sensors attractive for multirotor UAV applications. These sensors have low SWaP requirements and can record ¿16 M measurements at a time from the environment. These sensors can be combined with lens components that provide both wide-angle viewpoints, e.g., a 230 FOV via the fisheye lens, for omnidirectional perception and confined viewpoints, e.g., 80 FOV “standard” lens, for high fidelity target tracking and mapping. The intensive work required to integrate these sensors and develop optimized algorithms to process their data to work using onboard computing resources can be reused between daytime (EO) and nighttime (IR) sensing contexts.
Image sensors designed for both infrared (IR) and visible light often share common image processing algorithms, including basic processes like filtering, noise reduction, contrast enhancement, and image registration. Additionally, object detection, recognition, feature extraction, and image fusion algorithms can typically be adapted for both IR and visible light images, leveraging shared features and patterns. However, notable differences emerge, primarily related to spectral characteristics, illumination, noise, calibration, temperature considerations, environmental conditions, and the unique sensitivities of IR images to object materials. These distinctions necessitate adjustments in algorithms to address variations in contrast, object recognition, and material discrimination, showcasing the need for specialized approaches in certain contexts.
3.1.4 Recommendations
The assessment of available sensors suggests that the conventional image sensors are the best-practice sensors for multirotor UAV applications. EO and IR sensors, being lighter in weight and faster in measurement speeds compared to LiDAR sensors, also offer better image quality and nighttime measurement capability in contrast to event cameras. The choice of conventional sensors not only aligns with budgetary constraints but also caters to the diverse needs of UAV operations, encompassing navigation, mapping, and surveillance with exceptional performance and reliability.
@@ 3.2.1 Environment Simulation
AirSim [shah2018airsim] was used to simulate the dynamics of the drone. AirSim, developed by Microsoft, stands as a groundbreaking and influential simulator that has become a cornerstone in the development of autonomous drones and robotics. What sets AirSim apart is its capacity to simulate complex and dynamic environments with exceptional fidelity, replicating not only the physics of flight but also the intricacies of various sensors like cameras (RGB and depth), LiDAR, and GPS. Figure 5 shows an example of the AirSim simulated images captured by the cameras mounted on the multirotor.
a
b
c
d
Figure 5: An example of an AirSim city environment showing the following: (a) the FPV view in the simulator where the drone is hovering, (b) RGB image from the simulated left camera mounted on the drone, (c) RGB image from the simulated right camera, and (d) depth image from the simulated depth sensor where objects closer to the depth camera appear darker.
The proposed approach used the Cesium plugin for the Unreal Engine, also known as “Unreal Cesium”, to simulate real-world scenes, and enhance the effectiveness of simulations. Although AirSim provides rich virtual environments for testing and fine-tuning a wide array of autonomous systems, these environments are often designed for games and lack realism. To synthesize virtual models that replicate real-world contexts, AirSim can be integrated into the Unreal Engine to allow the Unreal Cesium plugin to create digital twins of real-world environment models. The Cesium plugin, given the latitude and longitude coordinates of desired locations, can load 3D tilesets at the location from Google Maps in the AirSim simulator.
The Unreal Cesium plugin creates a powerful combination by integrating the Unreal Engine’s advanced rendering and simulation capabilities with Cesium’s geo-spatial visualization and data streaming features. By streaming high-resolution 3D models from Google Maps, overlying them onto real-world maps, and applying dynamic lighting and shadows, to provide precise representations of real-world locations, such as cities, terrains, and 3D models of buildings. This integration allows developers to create highly realistic and spatially accurate virtual environments for various applications. Once the 3D map is generated, it behaves as a collision object in the Unreal Engine. The UAV then interacts with this model using the geometry of the environment and a physics engine. Figure 6 demonstrates the benefits of the Unreal Cesium plugin for simulation. It shows three real-world locations: (1) the UNC Charlotte campus, USA, (2) the Grand Canyon, USA, and (3) Paris, France.
a
b
c
Figure 6: Realistic environments created using Unreal Engine and Cesium Plug-in. (a) UNC Charlotte, NC, USA; (b) Grand Canyon, AZ, USA; (c) Eiffel Tower, Paris, France.
@@ 3.2.2 Flight Simulation
Figure 7 depicts the proposed pipeline for flight simulation. QGroundControl and PX4-Autopilot are software components commonly used in the field of UAVs and drones. They work together to provide a comprehensive solution for controlling and managing drone flights. The missions are planned by defining waypoints, flight paths, and specific actions for the drone to perform and QGroundControl sends the mission plans to the autopilot system PX4-Autopilot. As an open-source flight control software for UAVs, PX4-Autopilot runs on the flight controller onboard the drone and is responsible for stabilizing the aircraft, executing flight plans, and interfacing with sensors and actuators. Controlled by PX4-Autopilot, the simulated UAV in AirSim follows the planned trajectory in a high-realism virtual environment created by Unreal Engine and Cesium plug-in. The cameras mounted on the UAV then capture the images (RGB, depth, etc) of the scene. These images, along with the ground truth vehicle odometry, can be obtained from AirSim, which then can be applied together to generate the ground truth 3D model of the world.
Figure 7: The flight simulation pipeline integrates the following four robot development technologies to facilitate development and testing: (1) Unreal Engine and Cesium plugin (high-realism image synthesis), (2) AirSim (vehicle dynamics), (3) QGroundControl (mission planning), and (4) PX4-Autopilot (vehicle control and Software-In-The-Loop).
3.2.3 Ground Truth Geometry
The ground truth geometry is generated by applying AirSim’s built-in functionality for ground truth pose and noiseless telemetry to collect color-attributed point cloud data from simulated noiseless pixel-aligned RGB and depth images captured by a drone that traverses the environment. The telemetry is extracted from Cesium Tile data for generating high-fidelity geometric models. Poses and point clouds are integrated using standard mapping methods to reconstruct the scene geometry. Figure 8 shows an example of integrating the drone odometry (shown as the red curve Figure 8c) and the point cloud from RGB-D image sequences to create a geometric model of the scene.
@@ 3.3 Evaluation Methods
Mapping algorithm performance is evaluated using the following two key performance criteria: (1) mapping accuracy, and (2) mapping speed. Mapping accuracy is assessed by comparing the geometry of the reconstructed point cloud with the ground truth point cloud. Geometric accuracy measures each mapping algorithm’s ability to faithfully capture spatial relationships in the environment. Algorithm performance speed quantifies the amount of 3D estimates generated per unit of allocated computational resources. More computation allows more points to be tracked in sequential frames and the creation of more keyframes. Both tracked points and keyframe data feed non-linear bundle adjustment and batch trajectory optimization processes which improve map fidelity but can require significant computational resources. The evaluation methods allow result analysis that indicates design trade-offs associated with each mapping algorithm.
a
b
c
Figure 8: Ground truth geometry can be generated by transforming the point clouds of RGB-D frame sequences to the odometry of the drone which is shown in red in Figure 8(c). (a) An RGB frame captured by the drone; (b) A depth frame captured by the drone; (c) Integrating pose and point clouds to generate a map.
3.3.1 Point Cloud Registration
Point cloud registration seeks to compute the alignment between two 3D point clouds measured from the same surfaces in distinct coordinate systems. Alignment algorithms identify point correspondences between the misaligned point cloud datasets and compute the rigid Euclidean transformation that makes corresponding points coincide. To accomplish this, a point cloud is selected as the staticdataset and all other measured point clouds are transformed to align with the measurement coordinate system of the static dataset. Figure 9 shows an example of registering two point clouds where the red point cloud is the source set and the blue is the static dataset.
The Iterative Closest Point (ICP) algorithm [besl1992method] is employed to estimate point cloud alignments. The ICP algorithm consists of the following two steps: (1) compute correspondences and (2) compute the best alignment given the correspondence. Steps (1) and (2) are iterated until the alignment of Step (2) stops changing. Correspondences for a given iteration are calculated by finding the closest point in the static dataset to each point in the dataset being aligned. Closest point searches stop at a user-specified search radius for each point. The ICP algorithm seeks to minimize the RMSE (Root Mean Square Error) of all the distances between corresponding points and terminates when the gradient of RMSE is below a predefined threshold or a predetermined maximum iteration count is reached. Alignments resulting from the ICP algorithm are used to evaluate the geometry accuracy of mapping algorithms.
Figure 9: An example of point cloud registration [Zodage-2021-129203]. Red: source point cloud. Blue: target point cloud. Purple: registration result.